OpenStudy (anonymous):

f(x) = x+1/x+7 Find f^-1(-5)

OpenStudy (anonymous):

y = x+1/x+7 x/1 = y+1/y-3 x(y-3)=y+1 xy-3x=y+1 xy-y-3x=1 xy-y=3x+1 y(x-1)=3x+1 f^-1(x) = 3x+1/x-1 f^-1(5) = 3(-5)+1/-5-1=-15+1/-6=-16/-6=8/3

OpenStudy (anonymous):

I didn't get the correct answer

OpenStudy (anonymous):

@jim_thompson5910 @ganeshie8 @mathslover @jigglypuff314 @zzr0ck3r

jigglypuff314 (jigglypuff314):

y = x+1/x+7 x/1 = y+1/y-3 you changed the 7 to a 3?

OpenStudy (anonymous):

oh I didn't see that!

OpenStudy (anonymous):

that will change my answer

OpenStudy (anonymous):

I'm still trying so what I got was 17/3

jigglypuff314 (jigglypuff314):

x(y+7)=y+1 xy + 7x = y + 1 y - xy = 7x - 1 y (1-x) = 7x-1 y = (7x-1)/(1-x)

OpenStudy (anonymous):

where did the +1 change to a -1?

OpenStudy (anonymous):

@jigglypuff314

OpenStudy (anonymous):

you subtracted xy and -1 from both sides

OpenStudy (anonymous):

OK, I see what you did :)